OpenStudy (hippoharry):

The inverse of a function is represented as y^-1=x^2-2 . What is the equation of the original function?

OpenStudy (hippoharry):

\[a. y=-x2/2\] \[b. y=\sqrt{x^2-2}\] \[c. y=x^2+2\] \[d. y=\sqrt{x+2}\]

OpenStudy (hippoharry):

@satellite73

satellite73 (satellite73):

\[x=y^2-2\] solve for \(y\) in two steps

satellite73 (satellite73):

1) add 2 to both sides \[x+2=y^2\] then take the square root of both sides

satellite73 (satellite73):

got it?